OpenStudy (anonymous):

Add. (2c2 + 3c + 9) + (5c2 – 10c – 7)

OpenStudy (anonymous):

add the like terms Here i'll rearrange for you so hopefully you can see how to do it. (2c^2 + 5c^2) + (3c -10c) + (9-7)

OpenStudy (anonymous):

7c^2+3c-10c+9-7 7c+9-7 2 7c^2+7c+2? haha

OpenStudy (anonymous):

close. 3c-10c = -7c

OpenStudy (anonymous):

(2c2 + 3c + 9) + (5c2 – 10c – 7) take out of brackets = 2c^2+3c+9+5^2-10c-7 =2c^2+5c^2+3c-10c+9-7 you may want to put it in order, which makes it easier :) =7c^2-11c+2 add or subtract like terms

OpenStudy (anonymous):

@Ari_Wolff : the like terms for c dont equal -11c....again, 3c-10c = -7c.

OpenStudy (anonymous):

oops sorry @Puzzhang was doing a sum just like this one for homework. my bad. therefore the answer is: 7c^2-7c+2